# Incorporating Coupling Knowledge into Echo State Networks for Learning Spatiotemporally Chaotic Dynamics

### v1

@@ V Conclusion
Our work demonstrates that the coupling knowledge from an observed dynamical system can serve as a powerful guide to designing ESNs with a similarly coupled clustered reservoir layer. By conducting evaluations using benchmark dynamical systems, including the Lorenz-96 system and both homogeneous and inhomogeneous Kuramoto-Sivashinsky equations, we validate the effectiveness of our proposed PGC-ESN for both long-term attractor reconstruction and short-term prediction, even when prior coupling knowledge is slightly imperfect due to random rewiring with a low probability. By comparing PGC-ESN against a model with a clustered reservoir that lacks a coupling structure (RandC-ESN), we demonstrate that the effectiveness of PGC-ESN arises from the coupling structure in the reservoir. Additionally, by evaluating the models on the Lorenz-96 system with varied noise levels in the training data and systematic nonlinearity, we show that incorporating coupling knowledge into ESN models can also enhance their robustness to noise and enable superior predictive performance as the system’s nonlinearity increases.
@@
Beyond this study, we hope our work will inspire future research on integrating coupling knowledge and broader physical principles into the architectures of artificial neural networks and physical reservoir computers.Nakajima (2020) Physical systems are subject to various constraints, such as spatial limitations and input-channel restrictions, which affect the coupling between computational nodes. These constraints make it challenging to configure full connectivity and modulate existing connections. Therefore, designing computational architectures tailored to specific tasks is essential for physical computational systems. The findings of this study provide valuable insights into the relationship between information processing and computational mediums.
@@ Acknowledgements.
N. A. was supported by JSPS KAKENHI Grant No. JP22KJ1786 and JST PREST Grant No. JPMJPR24KF.
@@ A.1 Learning process for ESNs
A.2 Reservoir updation for physics-guided clustered ESNs
Fig. 7 provides a schematic illustration of the reservoir state update rule for PGC-ESNs when applied to a dynamical system in state $\bm{u}\in\mathbb{R}^{6}$ with a reservoir state $\bm{r}\in\mathbb{R}^{12}$. We focus on the updation of the $i^{\mathrm{th}}$ cluster in the reservoir state, which is highlighted in green. In both matrices $\bm{W}^{\mathrm{in}}$ (the input layer) and $\bm{W}$ (the reservoir layer), the symbol $\otimes$ indicates entries with a value of 0, and the blue grids represent the connections that transmit information from the coupled dimensions in $\bm{u}(t)$ or the coupled clusters in $\bm{r}(t)$. Meanwhile, the green grids transmit information from the corresponding dimension in $\bm{u}(t)$ or the cluster itself in $\bm{r}(t)$.
@@ A.2 Reservoir updation for physics-guided clustered ESNs
A.3 Physics-guided clustered ESNs for PDE target systems
Figure 8: Illustrative diagram of an example PDE system (a) and the learning process using the corresponding PGC-ESN (b).
Appendix B Numerical details on the evaluation metrics
B.1 Geometrical distance
The ideal invariant distributions $p(\bm{u})$ and $p(\bm{u}|\bm{\theta})$, which represent the state space densities of the ground truth and reconstructed systems, are approximated using Gaussian mixture models (GMMs) with the observed and predicted sequences $\{\bm{u}(1),\dots\bm{u}(T)\}$ and $\{\hat{\bm{u}}(1),\dots\hat{\bm{u}}(T)\}$ as follows:
@@ B.1 Geometrical distance
where each Gaussian distribution is centered at the observed or predicted state, respectively; the covariance matrix $\bm{\Sigma}$ determines the granularity of the spatial resolution. We set $\bm{\Sigma}=\mathbf{I}$ here, following the setting in Ref. Hess et al., 2023. The Kullback-Leibler (KL) divergence is then estimated using Monte Carlo approximation:
@@
where $n$ Monte Carlo samples $\bm{u}(i)$ are drawn from the GMM-approximated ground truth density $\hat{p}(\bm{u})$. To ensure that the measure is evaluated according to the system’s attractor rather than transient dynamics, we discard the first $25\%$ of time steps in both the observed and predicted sequences to obtain sequences of length $T$ for evaluation.
B.2 Temporal distance
[…]
Appendix C More evaluation results
@@
Figure 9: Evaluation results for the Lorenz-96 system under varied experimental settings. (a) Variation over training data size. (b) Variation over noise level $\kappa$. (c) Variation over $F$. (d) Variation over $\alpha$.
C.1 Lorenz-96 with varied experimental settings
Here, we present all evaluation results for the models under varied experimental settings with a reservoir size of 3000. First, we evaluate models tuned under standard settings but trained with different data sizes and noise levels $\kappa$. Specifically, we explore training data sizes ranging from $10^{5}$ down to $10^{4}$ in decrements of $10^{4}$ and noise levels from 0 to 200 in increments of 20.
Fig. 9 (a) presents the evaluation results across different training data sizes. Since Paral-ESN fails in the prediction phase when the training data size is smaller than $7\times 10^{4}$ in our experiment, we omit its results in (a). The findings reveal that our proposed PGC-ESN effectively reconstructs the system dynamics, maintaining $D_{\mathrm{geom}}$ around 14 and $D_{\mathrm{temp}}$ around 0.2, while achieving short-term prediction with VPT-1 around 2 Lyapunov times, even with a reduced training data size of $3\times 10^{4}$. In contrast, the predictive accuracy of the other two models deteriorates abruptly around this training data size. These results demonstrate that incorporating the coupling structure of the target system can enhance the ESN’s learning ability when the training data size is small.
Fig. 9 (b) presents the evaluation results under varying noise levels added to the training dataset. Similar to the $D_{\mathrm{temp}}$ plot described in Sec IV of the main paper, we can also see from the $D_{\mathrm{geom}}$ and VPT-1 plots that the performance of Paral-ESN and PGC-ESN declines more gradually than the other two models without incorporating the target system’s coupling structure. Additionally, our proposed PGC-ESN achieves the best short-term predictive power across the whole range of noise levels.
Fig. 9 (c) and (d) present all the evaluation results for variations in $F$ and $\alpha$ of the target system, respectively. In (c), we observe that similar to the $D_{\mathrm{temp}}$ plots described in Sec. IV of the main paper, the $D_{\mathrm{geom}}$ and VPT-1 plots show that our proposed PGC-ESN also achieves effective long-term attractor reconstruction and short-term prediction across the entire range of $F$, while the other models deteriorate as $F$ increases. In (d), it is evident that our proposed PGC-ESN excels in all three metric plots. Additionally, from the VPT-1 plot, we can see that PGC-ESN achieves effective short-term prediction up to $\alpha=6$, with over 3 Lyapunov times, while the other models deteriorate noticeably after $\alpha=3$.
@@ C.1 Lorenz-96 with varied experimental settings
Figure 10: Evaluation results for incorporating an imperfect coupling structure using the Lorenz-96 system.
C.2 Lorenz-96 with imperfect prior coupling knowledge
@@
Fig. 10 presents all the evaluation results for investigating the impact of using imperfect prior coupling knowledge in the Lorenz-96 system to guide the PGC-ESN’s reservoir structure. The plots for $D_{\mathrm{temp}}$ and VPT-1 are illustrated in Sec. IV of the main paper. Additionally, the $D_{\mathrm{geom}}$ plot shows the same trend that we discussed there.
C.3 Kuramoto–Sivashinsky
@@
Fig. 11 presents the reconstruction results for a specific target sequence of the homogeneous Kuramoto–Sivashinsky (KS) equation using a reservoir size of 2000 with the PGC-ESN, Paral-ESN, RandC-ESN, and Std-ESN (from left to right). Subplots (a) and (b) depict attractor reconstruction on the first two-dimensional subspace and power spectrum reconstructions, respectively. These results indicate that at a reservoir size of 2000, only PGC-ESN achieves effective long-term attractor reconstruction for this sequence. Subplot (c) further reveals that PGC-ESN provides the best short-term predictive power. Fig. 12 presents the results for the inhomogeneous KS equation under the same settings, indicating similar trends. In conclusion, our proposed PGC-ESN at reservoir size 2000 demonstrates strong capabilities for both long-term attractor reconstruction and short-term prediction with both the homogeneous and inhomogeneous KS equations.
@@
Figure 11: Predictive results for the homogeneous KS equation using a reservoir size of 2000 for PGC-ESN, Paral-ESN, RandC-ESN, and Std-ESN (from left to right). (a) The attractor on the first two-dimensional subspace. (b) The power spectrum. (c) Contour plots of the target sequence, the predicted sequence, and NRSE evolution across all dimensions, along with the NRMSE evolution averaged over all dimensions.
@@
Figure 12: Predictive results for the inhomogeneous KS equation using a reservoir size of 2000 for PGC-ESN, Paral-ESN, RandC-ESN, and Std-ESN (from left to right). (a) The attractor on the first two-dimensional subspace. (b) The power spectrum. (c) Contour plots of the target sequence, the predicted sequence, and NRSE evolution across all dimensions, along with the NRMSE evolution averaged over all dimensions.

### v2

@@ V Conclusion
Our work demonstrates that the coupling knowledge from an observed dynamical system can serve as a powerful guide to designing ESNs with a similarly coupled clustered reservoir layer. By conducting evaluations using benchmark dynamical systems, including the Lorenz-96 system and both homogeneous and inhomogeneous Kuramoto-Sivashinsky equations, we validate the effectiveness of our proposed PGC-ESN for both long-term attractor reconstruction and short-term prediction, even when the prior or extracted coupling knowledge is slightly imperfect. More specifically, incorporating coupling knowledge into ESNs leads to a substantial reduction in computational resources, such as reservoir size and the amount of training data. When resources are abundant, alternative approaches eventually match PGC-ESN in long-term attractor reconstruction but continue to underperform in short-term prediction. We attribute this trend to the inherent chaoticity of the underlying system, where sensitivity to initial conditions renders short-term prediction particularly challenging. In this context, incorporating prior coupling knowledge enables PGC-ESN to achieve superior state forecasting, even when computational resources are abundant. By contrast, the metrics used for long-term attractor reconstruction quantify the geometric and temporal discrepancies of the target and predicted sequences averaged over a prolonged time horizon. As a result, PGC-ESN’s advantage in this aspect diminishes once performance saturation is reached across all ESN models under sufficient computational resources.
By comparing PGC-ESN against a model with a clustered reservoir that lacks a coupling structure (RandC-ESN), we demonstrate that the effectiveness of PGC-ESN arises from the coupling structure in the reservoir. Additionally, by evaluating the models on the Lorenz-96 system with varied experimental settings, we show that compared to Std-ESN and RandC-ESN—which do not utilize coupling knowledge—incorporating coupling knowledge into ESN models, as done in PGC-ESN and Paral-ESN, enhances their robustness to variations in training noise level, training data size, and sequence temporal resolution, while also enabling superior predictive performance as the target system’s nonlinearity increases.
@@
However, our method has a restriction on computational efficiency compared to Paral-ESN. Paral-ESN is known for its superior scalability, owing to its modular architecture and the possibility of sequential training, which makes it particularly well-suited for learning large-scale spatiotemporal systems, such as 2- or 3-dimensional PDE systems. For homogeneous systems, training a single reservoir and replicating it spatially is often sufficient when using Paral-ESN. Vlachas et al. (2020) Given the computational efficiency of Paral-ESN and prior work incorporating fully coupled structures into Paral-ESN for learning chaotic systems, Baur and Räth (2021) we believe that integrating the system’s coupling knowledge into Paral-ESN’s framework represents a promising direction for future research. Moreover, by leveraging spatial translation symmetry, a trained Paral-ESN can generalize to systems with varying spatial extent. Goldmann et al. (2022) Prior studies have also demonstrated Paral-ESN’s flexibility to integrate with dimensionality reduction techniques. Fleddermann, Parlitz, and Wellecke (2025) These insights suggest that incorporating system-specific knowledge, such as coupling structure, into ESN models in conjunction with the aforementioned and more general techniques for ESNs holds significant potential for advancing future research.
Our results also suggest a future direction in which the reservoir could self-organize to optimize its network configuration. In this study, we showed that even without prior knowledge of the coupling structure, incorporating coupling information extracted via external time-series analysis enhances ESN’s predictive performance. Banerjee et al. Banerjee et al. (2021) demonstrated that a reservoir model can perform link inference for dynamics directly from time series data. This implies that the reservoir itself could simultaneously extract coupling knowledge from time series data and use it to self-organize its internal structure accordingly. Furthermore, the observed contribution of the similarity between the reservoir and target system structures for predictive performance may help explain why physical reservoir computing Nakajima (2020) excels at predicting and controlling the dynamics of the physical system itself. Goto, Nakajima, and Notsu (2021); Akashi et al. (2024)
Beyond this study, we hope our work will inspire future research on integrating coupling knowledge and broader physical principles into the architectures of artificial neural networks and physical reservoir computers. Physical systems are subject to various constraints, such as spatial limitations and input-channel restrictions, which affect the coupling between computational nodes. These constraints make it challenging to configure full connectivity and modulate existing connections. Therefore, designing computational architectures tailored to specific tasks is essential for physical computational systems. The findings of this study provide valuable insights into the relationship between information processing and computational medium.
@@ Acknowledgements.
N. A. was supported by JSPS KAKENHI Grant No. JP22KJ1786, 25K00011, and JST PREST Grant No. JPMJPR24KF.
@@ A.1 Learning process for ESNs
A.2 Physics-guided clustered ESNs for PDE target systems
Figure 8: Illustrative diagram of an example target PDE system (a) and the learning process using the corresponding PGC-ESN (b).
A.3 Reservoir updation for physics-guided clustered ESNs
Fig. 9 provides a schematic illustration of the reservoir state update rule for PGC-ESNs with a reservoir state $\bm{r}\in\mathbb{R}^{12}$ when applied to a dynamical system with state $\bm{u}\in\mathbb{R}^{6}$. We focus on the updation of the $i^{\mathrm{th}}$ cluster in the reservoir state. The $i^{\mathrm{th}}$ dimension of the input state, along with the corresponding cluster in the reservoir, are highlighted in green, while the coupled dimensions and clusters are shown in blue. In both matrices $\bm{W}^{\mathrm{in}}$ (the input layer) and $\bm{W}$ (the reservoir layer), the symbol $\otimes$ indicates entries with a value of 0, and the blue grids represent the connections that transmit information from the coupled dimensions in $\bm{u}(t)$ or the coupled clusters in $\bm{r}(t-1)$. Meanwhile, the green grids transmit information from the corresponding dimension in $\bm{u}(t)$ or the cluster itself in $\bm{r}(t-1)$.
@@ A.3 Reservoir updation for physics-guided clustered ESNs
Appendix B Definition and numerical details on transfer entropy
Transfer Entropy (TE) is an information-theoretic measure that quantifies the directed information flow between two time series. Schreiber (2000) Let $X_{t}$ and $Y_{t}$ for $t\in\mathbb{N}$ be two stochastic processes, the transfer entropy from $Y$ to $X$ is defined as the conditional mutual information between the next state $X_{t+1}$ and the past of $Y$, conditioned on the past of $X$:
$T_{Y\to X}$ $=I(X_{t+1};\bm{Y}_{t}^{(L)}\mid\bm{X}_{t}^{(L)})$
$=\iiint p(x_{t+1},\bm{x}_{t}^{(L)},\bm{y}_{t}^{(L)})\log\left(\frac{p(x_{t+1}\mid\bm{x}_{t}^{(L)},\bm{y}_{t}^{(L)})}{p(x_{t+1}\mid\bm{x}_{t}^{(L)})}\right)\,$
$\quad\quad\quad dx_{t+1}\,d\bm{x}_{t}^{(L)}\,d\bm{y}_{t}^{(L)}$ (14)
where $I(\cdot\,;\,\cdot\mid\cdot)$ denotes the conditional mutual information, $\bm{X}_{t}^{(L)}=[X_{t-L+1},\dots,X_{t-1},X_{t}]$ is the length-$L$ past state vector of $X$, and $\bm{Y}_{t}^{(L)}=[Y_{t-L+1},\dots,Y_{t-1},Y_{t}]$ is the length-$L$ past state vector of $Y$.
In our implementation, the calculation of transfer entropy is restricted to pairs of univariate time series, as it is used solely to infer the coupling strengths between individual dimensions of the Lorenz-96 system in this study. We set the length of the past state vectors to $L=5$ and retain only the earliest state in each vector. To approximate the underlying probability distributions, each time series is discretized using histogram binning, with the number of bins automatically determined via the Freedman–Diaconis rule. Freedman and Diaconis (1981) To reduce estimation noise, the approximated probability distributions are smoothed using a Gaussian filter with standard deviation $\sigma=1$. The final integration is approximated by summing the smoothed probabilities over all bins.
Appendix C Numerical details on the evaluation metrics
C.1 Geometrical distance
To ensure that the measure is evaluated according to the system’s attractor rather than transient dynamics, we discard the first $25\%$ of time steps in both the observed and predicted sequences to obtain sequences of length $T$ for evaluation. The ideal invariant distributions $p(\bm{u})$ and $p(\bm{u}|\bm{\theta})$, which represent the state space densities of the ground truth and reconstructed systems, are approximated via Gaussian mixture models (GMMs) using the observed and predicted sequences $\{\bm{u}(1),\dots\bm{u}(T)\}$ and $\{\hat{\bm{u}}(1),\dots\hat{\bm{u}}(T)\}$, respectively:
@@ C.1 Geometrical distance
where each Gaussian distribution is centered at an observed or predicted state, respectively; the covariance matrix $\bm{\Sigma}$ determines the granularity of the spatial resolution. We set $\bm{\Sigma}=\mathbf{I}$ here, following the setting in Ref. Hess et al., 2023. The Kullback-Leibler (KL) divergence is then estimated using a Monte Carlo approximation:
@@
where $n=1000$ Monte Carlo samples $\bm{u}(i)$ are drawn individually from a randomly selected individual Gaussian component of the GMM approximating $\hat{p}(\bm{u})$.
@@
Figure 10: Evaluation results for the Lorenz-96 system under varied experimental settings. (a) Variation over training data size. (b) Variation over noise level $\kappa$. (c) Variation over sequence temporal resolution. (d) Variation over the target system’s forcing term $F$. (e) Variation over the target system’s nonlinearity parameter $\alpha$.
C.2 Temporal distance
As with the geometrical distance $D_{\mathrm{geom}}$, we discard the first $25\%$ of time steps in the observed and predicted sequences to eliminate the influence of transient dynamics. We then standardize each dimension of both sequences $\{\bm{u}(1),\dots\bm{u}(T)\}$ and $\{\hat{\bm{u}}(1),\dots\hat{\bm{u}}(T)\}$ so that $f_{i}$ and $g_{i}$ satisfy $\int_{-\infty}^{\infty}f_{i}(\omega)\,d\omega=1$ and $\int_{-\infty}^{\infty}g_{i}(\omega)\,d\omega=1$, which is required to compute the Hellinger distance $H$. We approximate the power spectra $f_{i}$ and $g_{i}$ using a Fast Fourier Transform (FFT) Cooley and Tukey (1965), yielding $\hat{\bm{f}}_{i}=|\mathcal{F}u_{i,1:T}|^{2}$ and $\hat{\bm{g}}_{i}=|\mathcal{F}\hat{u}_{i,1:T}|^{2}$, where $\hat{\bm{f}}_{i}$ and $\hat{\bm{g}}_{i}$ are the discrete power spectra for the $i^{\mathrm{th}}$ dimension. The approximated power spectra are then smoothed using a Gaussian kernel with a standard deviation $\sigma=2$ and a window length $l=10\sigma+1$. The high-frequency tails beyond the first 500 frequency components are cut off. Finally, the trapezoidal rule is applied to numerically integrate the pointwise geometric mean of the discrete power spectra over frequency.
@@ C.2 Temporal distance
Figure 11: Evaluation results for incorporating imperfect prior coupling knowledge using the Lorenz-96 system.
Appendix D More evaluation results
D.1 Lorenz-96 with varied experimental settings
Here, we present all evaluation results for the models under varied experimental settings with a reservoir size of 3000. First, we evaluate models tuned under standard settings but trained with different data sizes, noise levels $\kappa$, and sequence temporal resolution.
Fig. 10 (a) presents the evaluation results across different training data sizes, ranging from $10^{5}$ down to $10^{4}$ in decrements of $10^{4}$. The findings reveal that our proposed PGC-ESN effectively reconstructs the system dynamics, maintaining $D_{\mathrm{geom}}$ around 14 and $D_{\mathrm{temp}}$ around 0.12, while achieving short-term prediction with VPT-1 around 2 Lyapunov times, even with a reduced training data size of $10^{4}$. Paral-ESN also exhibits moderate sensitivity to reductions in training data size, achieving the lowest $D_{\mathrm{geom}}$ values across the entire range, while maintaining inferior performance in $D_{\mathrm{temp}}$ and VPT-1 with mild deterioration along the data reduction. In contrast, the predictive accuracy of the other two models without incorporating coupling knowledge deteriorates abruptly when the training data size falls below $3\times 10^{4}$. These results demonstrate that incorporating the coupling structure of the target system—as done in PGC-ESN and Paral-ESN—can enhance the ESN’s learning ability when the training data size is small.
Fig. 10 (b) presents the evaluation results under varying noise levels added to the training dataset. Similar to the $D_{\mathrm{temp}}$ plot described in Sec. IV E of the main paper, we can also see from the $D_{\mathrm{geom}}$ and VPT-1 plots that the performance of Paral-ESN and PGC-ESN declines more gradually than the other two models without incorporating the target system’s coupling structure. Additionally, our proposed PGC-ESN achieves the best short-term predictive power across the whole range of noise levels.
Fig. 10 (c) presents the evaluation results under varying sequence temporal resolutions. Consistent with the $D_{\mathrm{temp}}$ plot discussed in Sec. IV E of the main paper, the $D_{\mathrm{geom}}$ results indicate that PGC-ESN and Paral-ESN exhibit a more gradual decline in long-term attractor reconstruction performance as the step size increased, compared to Std-ESN and RandC-ESN. In contrast, short-term prediction performance deteriorates at a similar rate across all models. Note that when the step size reaches $5\Delta t$, the $D_{\mathrm{geom}}$ values for all models diverge to infinity and are therefore omitted from the plot.
Next, we tune and evaluate all models on the Lorenz-96 system under varying values of the forcing term $F$ and the nonlinearity parameter $\alpha$. Fig. 10(d) and (e) present all the evaluation results for variations in $F$ and $\alpha$ of the target system, respectively. In (d), we observe that similar to the $D_{\mathrm{temp}}$ plots described in Sec. IV E of the main paper, the $D_{\mathrm{geom}}$ and VPT-1 plots show that our proposed PGC-ESN also achieves effective long-term attractor reconstruction and short-term prediction across the entire range of $F$, while the other models deteriorate as $F$ increases. In (e), it is evident that our proposed PGC-ESN excels in all three metric plots. Additionally, from the VPT-1 plot, we can see that PGC-ESN achieves effective short-term prediction up to $\alpha=6$, with over 3 Lyapunov times, while the other models deteriorate noticeably after $\alpha=3$.
D.2 Lorenz-96 with imperfect prior coupling knowledge
@@
Fig. 11 presents all the evaluation results for investigating the impact of using imperfect prior coupling knowledge in the Lorenz-96 system to guide the PGC-ESN’s reservoir structure. The plots for $D_{\mathrm{temp}}$ and VPT-1 are illustrated in Sec. IV F of the main paper. Additionally, the $D_{\mathrm{geom}}$ plot shows the same trend that we discussed there.
D.3 Kuramoto–Sivashinsky
@@
Fig. 12 presents the reconstruction results for a specific target sequence of the homogeneous Kuramoto–Sivashinsky (KS) equation using a reservoir size of 2000 with the PGC-ESN, Paral-ESN, RandC-ESN, and Std-ESN (from left to right). Subplots (a) and (b) depict attractor reconstruction on the first three-dimensional subspace and power spectrum reconstructions, respectively. These results indicate that at a reservoir size of 2000, PGC-ESN achieves the best long-term attractor reconstruction for this sequence. Subplot (c) further reveals that PGC-ESN provides the best short-term predictive power. Fig. 13 presents the results for the inhomogeneous KS equation under the same settings, indicating similar findings. In conclusion, our proposed PGC-ESN at reservoir size 2000 demonstrates strong capabilities for both long-term attractor reconstruction and short-term prediction with both the homogeneous and inhomogeneous KS equations.
@@
Figure 12: Predictive results for the homogeneous KS equation using a reservoir size of 2000 for PGC-ESN, Paral-ESN, RandC-ESN, and Std-ESN (from left to right). (a) The attractor on the first three-dimensional subspace. (b) The power spectrum. (c) Contour plots of the target and predicted sequences (d) NRMSE evolution and contour plots of the NRSE evolution.
@@
Figure 13: Predictive results for the inhomogeneous KS equation using a reservoir size of 2000 for PGC-ESN, Paral-ESN, RandC-ESN, and Std-ESN (from left to right). (a) The attractor on the first three-dimensional subspace. (b) The power spectrum. (c) Contour plots of the target and predicted sequences (d) NRMSE evolution and contour plots of the NRSE evolution.